OpenStudy (kitkat16):

Please Help!!! :( Carlos is analyzing the results of an experiment on two groups of mice in which he compared the difference in their weight gain after feeding each group a different type of food for one month. The difference in the mean weight gain is 0.6 grams. The standard deviation of the difference in sample means is 0.305. The 68% confidence interval for the population mean difference is ±??? The 95% confidence interval for the population mean difference is ±???

OpenStudy (kitkat16):

so do I first have to look for the z score? z=-0.51

OpenStudy (perl):

Confidence interval for population mean difference = difference in sample mean weight ± z score * standard deviation

OpenStudy (perl):

we know that , by the empirical rule, within one standard deviation is 68% of the data and 95% of the data is within 2 standard deviations. Therefore the z score for 68% two tailed is z=1, and z score for 95% two tailed is z = 2. 68% confidence interval: .6 ± 1* 0.305 95% confidence interval: .6 ± 2* 0.305

OpenStudy (kitkat16):

so would I write it like this? The 68% confidence interval for the population mean difference is ± 0.905 The 95% confidence interval for the population mean difference is ± 1.21

OpenStudy (perl):

not quite

OpenStudy (perl):

.6 ± 1* 0.305 = ( .6 - 0.305 , .6 + 0.305)

OpenStudy (kitkat16):

for this one it isn't correct. The 95% confidence interval for the population mean difference is ± 1.21 it is one of these 0.160 , 0.305, 0.610, 0.915

OpenStudy (perl):

.6 ± 1* 0.305 = ( .6 - 0.305 , .6 + 0.305) = ( .295, .905) .6 ± 2* 0.305 = ( -.01, 1.21)

OpenStudy (kitkat16):

They would both be 0.305

OpenStudy (perl):

can you take a screen shot of this?

OpenStudy (perl):

I think it should say .6 ± .305, but for some reason they left out the .6

OpenStudy (kitkat16):

Im confused as to why they would be the same .305

OpenStudy (perl):

because it can be greater or less than .305

OpenStudy (perl):